# Unveiling User Preferences: A Knowledge Graph and LLM-Driven Approach for Conversational Recommendation

### v1

@@ 3.2.2. Graph-to-Text Adapter
3.2.3. Large Language Model (LLM)
The LLM is the main reasoning engine that generates user preferences by synthesizing information from dialogue histories and KG-augmented contexts. It captures both the explicit user preferences expressed in the dialogue and the implicit preferences inferred from the KG-augmented context for accurate preference analysis. Our framework is compatible with various state-of-the-art LLMs, allowing flexibility in model choice. In this paper, we use Llama3.1-8B (Dubey et al., 2024) for its natural language understanding and generation capabilities. To adapt the chosen LLM efficiently for our task, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) for fine-tuning.
3.3. Training Pipeline
3.3.1. Graph Entity Captioning
To obtain a Graph-to-Text adapter that effectively bridges the modality gap between the graph entity embeddings and the LLM’s semantic space, inspired by the pre-training strategies employed in vision-language models (Dai et al., 2024; Liu et al., 2023a), we introduce a graph entity captioning pretraining mechanism, as shown in Figure 2a. This process creates a strong connection between graph-structured data and natural language by generating entity-specific captions that encompass both the entity’s intrinsic information and aggregated data from its neighboring nodes, simulating a message-passing process within the graph structure.
The pre-training stage enables the LLM to interpret graph entity embeddings in a semantic context, leading to a deeper understanding of the relationships and attributes encoded within the KG, which is crucial for subsequent preference modeling. The caption generation process differentiates between item entities (e.g., movies) and non-item entities (e.g., genres, directors). For item entities, we employ a structured template that captures key attributes and relationships.
Example of Item Entity Caption Template Movie Title: ¡Title¿; Genres: ¡Genres¿; Keywords: ¡Keywords¿; Starring: ¡Actors¿; Directors: ¡Directors¿; Production Company: ¡Company¿; Plot: ¡Plot Summary¿
This template ensures comprehensive coverage of item attributes while maintaining a consistent structure across different items. For non-item entities, we adopt a more flexible approach that emphasizes the entity’s role and its connections within the KG  [^1] Details on caption construction and instruction templates for both item and non-item entities are provided in Appendix B.2..
@@ 3.3.1. Graph Entity Captioning
Entity Captioning Input: ¡Entity Embeddings $\mathbf{h}_{e}^{\tau}$, ¡Instruction $I_{c}$¿ Output: ¡Entity Caption $C_{e}$¿
@@
(6) $\mathcal{L}_{caption}=-\sum_{e\in\mathcal{E}}\log P(C_{e}\mid\mathbf{h}_{e}^{\tau},I_{c}),$
@@
3.3.2. Knowledge-aware Instruction Tuning
After the graph entity captioning pre-training, the LLM has gained a basic understanding of the KG structure and content. However, it has not yet been explicitly trained to utilize this knowledge for downstream tasks such as preference modeling and recommendation generation. To this end, we introduce a knowledge-aware instruction tuning, which aims to enable the LLM with abilities of reasoning across modalities, integrating information from both the dialogue history and the KG to infer user preferences and interests, as shown in Figure 2b.
@@ 3.3.2. Knowledge-aware Instruction Tuning
This prompt follows a coarse-to-fine structure containing four key steps: (1) reasoning, providing transparency in the model’s decision-making; (2) overall preferences, offering a broad view of the user’s tastes; (3) current interests, capturing recent and specific preferences to guide subsequent recommendations; and (4) recommendation, leveraging the LLM’s reasoning capabilities to suggest relevant items aligned with user preferences, guiding downstream CRS models. This structured prompt ensures that the model captures both long-term preferences and immediate interests. To generate ground-truth preference summaries, we utilize an advanced LLM (e.g., ChatGPT) that performs cross-modal reasoning, integrating complete dialogue histories with structured metadata of mentioned items from the KG. The detailed process and instruction templates for this ground-truth generation are provided in Appendix B.3.
The instruction tuning process can be represented in the following format:
Knowledge-aware Instruction Tuning Input: ¡Mentioned Entities Embeddings $\mathbf{E}_{t}$¿, ¡Instruction $I_{p}$¿, ¡Dialogue History $H_{t}$¿ Output: ¡Preference Summary $\mathcal{P}_{t}$¿
A detailed example of the preference summary can be found in Table 5 of Appendix C.2.
This instruction tuning process is optimized by minimizing the NLL of the generated preference summaries, as expressed by:
(7) $\mathcal{L}_{preference}=-\sum_{\mathcal{D}\in\mathcal{C}}\sum_{H_{t}\in\mathcal{D}}\log P(\mathcal{P}_{t}\mid\mathbf{E}_{t},I_{p},H_{t}),$
@@
3.4. Integration with Existing CRS Models
@@
3.4.1. Preference Representation
Traditional CRSs are not designed to directly utilize natural language preference summaries. Therefore, we explore two methods to encode these summaries into a format suitable for existing CRS architectures:
1) PLM-based representation: This method leverages a PLM to extract rich semantic representations from text. The PLM is adaptable and can be implemented as either a frozen or trainable text encoding model. In our implementation, we employ BERT (Devlin et al., 2019) to encode the user preference summary as follows:
(8) $\bm{s}_{c}^{\text{text}}=\text{PLM}(\mathcal{P}),$
where $\bm{s}_{c}^{\text{text}}$ is the encoded preference from text preference $\mathcal{P}$, specifically the [CLS] token embedding. The contextual understanding of language models enables $\bm{s}_{c}^{\text{text}}$ to capture comprehensive user preference summaries at both coarse-grained (i.e., overall user preferences) and fine-grained levels (i.e., current interests and specific items).
2) EOS representation: Considering the auto-regressive nature of LLMs and the information-rich preference summaries generated by COMPASS, we implement a computationally efficient encoding method. This approach leverages the [EOS] token embedding from the COMPASS-generated summary as it encapsulates the cumulative context of the entire preference description. We then process this embedding through a lightweight two-layer Multi-Layer Perceptron (MLP) to encode the preference as follows:
(9) $\bm{s}_{c}^{\text{EOS}}=\text{MLP}(\bm{z}_{\text{EOS}}),$
where $\bm{z}_{\text{EOS}}$ is the [EOS] token embedding from the COMPASS-generated preference summary, and $\bm{s}_{c}^{\text{EOS}}$ is the resulting encoded preference. This approach offers computational efficiency while still capturing the essential information from the preference summary, enabling effective integration with existing CRS models. To streamline the notation, we simplify $\bm{s}_{c}^{\text{text}}$ and $\bm{s}_{c}^{\text{EOS}}$ to $\bm{s}_{c}$.
3.4.2. Enhanced Recommendation
We integrate the encoded preferences into the existing CRS models to enhance their recommendation performance. We employ an adaptive gating mechanism to enhance the preference representation $\bm{s}_{b}$ captured from the base CRS model [^2] The preference representation $\bm{s}_{b}$ is specific to each CRS model as different models employ distinct methods for building user preferences. with our COMPASS-generated representation $\bm{s}_{c}$. Formally, we have:
(10) $\gamma$ $=\sigma(\mathbf{W}[\bm{s}_{b};\bm{s}_{c}]),$
(11) $\bm{s}_{u}$ $=\gamma\bm{s}_{b}+(1-\gamma)\bm{s}_{c},$
@@ 3.4.2. Enhanced Recommendation
(12) $P_{rec}(i)=\text{softmax}(\bm{s}_{u}\cdot\bm{I}_{i}),$
@@
(13) $\mathcal{L}_{\text{rec}}=-\sum_{j=1}^{N}\sum_{i=1}^{M}y_{ij}\log\left(\bm{P}_{\text{rec}}^{(j)}(i)\right),$
@@
4. Experiments
4.1. Experimental Settings
4.1.1. Datasets
We conduct our experiments on two widely used English CRS datasets: ReDial (Li et al., 2018) and INSPIRED (Hayati et al., 2020) datasets. The ReDial dataset, focused on movie recommendations, contains 11,348 conversations and is constructed through crowdsourcing on Amazon Mechanical Turk (AMT). The INSPIRED dataset contains 999 conversations, also about movie recommendations, but additionally provides recommendation strategies based on social science theories. We constructed the knowledge graph by scraping data from IMDB [^4] https://www.imdb.com/, using movie titles and release years as key search terms.
4.1.2. Baselines
We consider a comprehensive range of baseline models, including traditional CRS ReDial (Li et al., 2018), knowledge-graph based methods such as KBRD (Chen et al., 2019) and KGSF (Zhou et al., 2020), language model-based approaches like BERT (Devlin et al., 2019), GPT-2 (Radford et al., 2019), Llama3.1-8B (Dubey et al., 2024), and hybrid models combining KGs with language models such as BARCOR (Wang et al., 2022b) and PECRS (Ravaut et al., 2024).
Since no existing baselines specifically generate user preference summaries, we also compare COMPASS to Llama3.1-8B and GPT-4o, with Llama3.1-8B denoted as Llama-Summary when used solely for generating user preference summaries without access to the KG. More details on these models can be found in Appendix B.1.
4.1.3. Evaluation Metrics
Our evaluation framework assesses both the performance of the recommendations and the quality of generated user preference summaries. For recommendation tasks, we adopt widely used metrics from previous works (Chen and Sun, 2023; Zhou et al., 2022, 2020), including HR@$K$, NDCG@$K$, and MRR@$K$, with $K$ set to 10 and 50. To evaluate user preference summaries, we employ three types of metrics: (1) Lexical Similarity: ROUGE-1, ROUGE-2, ROUGE-L, and ROUGE-Sum; (2) Semantic Understanding: Following (Wang et al., 2023a), we use GPT-4o-mini to assess Reasoning Proficiency (RP) and Factual Consistency (FC); (3) User-Centric Evaluation: LLM-based simulated user evaluations, inspired by (Ji et al., 2023; Gao et al., 2024), measure Explainability and User Preference Alignment. More details on the metrics used can be found in Appendix B.4.
4.1.4. Implementation Details
We implement COMPASS using the Llama3.1-8B model (Dubey et al., 2024) as the base LLM, which consists of 32 transformer layers with an embedding dimension of 4096. We freeze all parameters of the base LLM and employ LoRA (Hu et al., 2022) for fine-tuning. For the graph encoder, we set the number of layers to 1, with a hidden dimension of 768. We use batch sizes of 256 for pre-training and 128 for fine-tuning on the Inspired dataset. For the ReDial dataset, we maintain a batch size of 128 for both pre-training and fine-tuning. Early stopping is implemented to optimize training. All experiments are conducted on Nvidia A100 GPUs. For ground truth user preference summaries, we utilize the OpenAI API. More details on this process are provided in Appendix B.3. Our code will be made publicly available upon acceptance.
Table 1. Performance comparison on recommendation tasks. ‘COM.’ denotes models enhanced with the COMPASS approach and ‘Improv.” indicates the relative improvement of the COMPASS compared to the original base model. The best results are highlighted in bold
@@ 4.1.4. Implementation Details
Base BERT (Devlin et al., 2019) 0.143 0.319 0.073 0.108 0.052 0.059 0.179 0.328 0.095 0.125 0.072 0.079
GPT-2 (Radford et al., 2019) 0.147 0.327 0.071 0.107 0.051 0.056 0.112 0.278 0.089 0.128 0.063 0.076
Llama3.1-8B (Dubey et al., 2024) 0.188 0.376 0.103 0.146 0.078 0.087 0.190 0.332 0.118 0.150 0.094 0.102
ReDial (Li et al., 2018) 0.140 0.320 0.061 0.065 0.035 0.045 0.117 0.285 0.035 0.072 0.022 0.048
KBRD (Chen et al., 2019) 0.151 0.336 0.099 0.136 0.071 0.079 0.172 0.265 0.106 0.127 0.086 0.091
KGSF (Zhou et al., 2020) 0.183 0.378 0.098 0.140 0.072 0.081 0.175 0.273 0.106 0.128 0.088 0.093
BARCOR (Wang et al., 2022b) 0.169 0.374 0.088 0.133 0.063 0.073 0.185 0.339 0.104 0.137 0.080 0.087
PECRS (Ravaut et al., 2024) 0.205 0.399 0.112 0.154 0.083 0.093 0.179 0.337 0.106 0.142 0.084 0.092
@@
4.2. Recommendation Evaluation
4.2.1. Improvement over Baseline Models
@@
COMPASS is designed to be flexible, allowing integration with different CRS models. To assess its effectiveness, we evaluated COMPASS across various CRS models. Performance results are shown in Table 1. We have the following observations: (1) Baseline models incorporating external KGs, such as KBRD and KGSF, consistently outperform simpler language model-based approaches like BERT and GPT-2. This highlights the importance of structured knowledge in capturing user preferences and item relationships, particularly in conversational recommendation settings. (2) Methods that combine language models with KGs, such as BARCOR and PECRS, show further improvements over KG-only models, demonstrating the benefits of integrating both sources of information. (3) Llama3.1-8B demonstrates strong performance, surpassing BARCOR and performing competitively with PECRS on the ReDial dataset while outperforming it on the INSPIRED dataset. This indicates that the advanced language understanding and extensive world knowledge of LLMs can effectively capture accurate user preferences, resulting in more accurate recommendations.
@@
Building upon these strong baseline performances, the integration of COMPASS with these models leads to substantial improvements across all evaluation metrics. Notably, when integrated with KBRD on the INSPIRED dataset, COMPASS achieves remarkable relative improvements of 44.76% in HR@10 and 47.92% in HR@50, outperforming all baseline models, including PECRS. When applied to the already strong Llama3-8B model, COMPASS still delivers considerable enhancements, with increases of 13.76% in HR@10 and 14.45% in NDCG@10 on the ReDial dataset, and even greater gains on the INSPIRED dataset. These consistent improvements are attributed to the high-quality, knowledge-enriched user preference representations generated by COMPASS. These representations not only capture user intent more effectively but also provide structured insights that enhance the overall recommendation process, leading to both better performance and explainability.
4.2.2. Comparison of Enhancement Methods

### v2

@@ IV-B2 Graph-to-Text Adapter
IV-B3 Large Language Model (LLM)
The LLM is the main reasoning engine that generates user preferences by synthesizing information from dialogue histories and KG-augmented contexts. It captures both the explicit user preferences expressed in the dialogue and the implicit preferences inferred from the KG-augmented context for accurate preference reasoning. Our framework is compatible with various state-of-the-art LLMs, allowing flexibility in model choice.
IV-C Training Pipeline
IV-C1 Graph Entity Captioning
To obtain a Graph-to-Text adapter that bridges the modality gap between the graph entity embeddings and the LLM’s semantic space, inspired by the pre-training strategies employed in vision-language models [33, 34], we introduce a graph entity captioning pre-training mechanism, as shown in Figure 2a. This process creates a strong connection between graph-structured data and natural language by generating entity-specific captions that combine intrinsic entity information with neighboring node data, simulating message-passing within the graph structure. The pre-training stage enables the LLM to interpret graph entity embeddings in a semantic context, facilitating deeper understanding of KG relationships and attributes for preference modeling.
The caption generation process differentiates between item entities (e.g., movies) and non-item entities (e.g., genres, directors). For item entities, we employ a structured template that captures key attributes and relationships.
Example of Item Entity Caption Template Movie Title: <Title>; Genres: <Genres>; Keywords: <Keywords>; Starring: <Actors>; Directors: <Directors>; Production Company: <Company>; Plot: <Plot Summary>
This template ensures comprehensive coverage of item attributes while maintaining a consistent structure across different items. For non-item entities, we adopt a more flexible approach that emphasizes the entity’s role and its connections within the KG.
@@ IV-C1 Graph Entity Captioning
Entity Captioning Input: <Entity Embeddings $\mathbf{h}_{e}^{\tau}$ >, <Instruction $I_{c}$>
Output: <Entity Caption $C_{e}$>
@@
$\mathcal{L}_{caption}=-\sum_{e\in\mathcal{E}}\log P(C_{e}\mid\mathbf{h}_{e}^{\tau},I_{c}),$ (6)
@@
IV-C2 Knowledge-aware Instruction Tuning
After the graph entity captioning pre-training, the LLM has gained a basic understanding of the KG structure and content. However, it has not yet been explicitly trained to utilize this knowledge for downstream tasks such as preference modeling and recommendation generation. To this end, we introduce knowledge-aware instruction tuning to enable the LLM to reason across modalities, integrating dialogue history and KG information to infer user preferences and interests, as shown in Figure 2b.
@@ IV-C2 Knowledge-aware Instruction Tuning
This prompt follows a coarse-to-fine structure containing four key steps: (1) reasoning, providing transparency in the model’s decision-making; (2) overall preferences, offering a broad view of the user’s tastes; (3) current interests, capturing recent and specific preferences to guide subsequent recommendations; and (4) recommendation, leveraging the LLM’s reasoning capabilities to suggest relevant items aligned with user preferences, guiding downstream CRS models. This structured prompt ensures that the model captures both long-term preferences and immediate interests. To generate ground-truth preference summaries, we utilize an advanced LLM (e.g., ChatGPT) that performs cross-modal reasoning, integrating complete dialogue histories with structured metadata of mentioned items from the KG. The instruction tuning process can be represented in the following format:
Knowledge-aware Instruction Tuning Input: <Mentioned Entities Embeddings $\mathbf{E}_{t}$>, <Instruction $I_{p}$>, <Dialogue History $H_{t}$>
Output: <Preference Summary $\mathcal{P}_{t}$>
A detailed example of the preference summary can be found in Table IV. This instruction tuning process is optimized by minimizing the NLL of the generated preference summaries, as expressed by:
$\mathcal{L}_{preference}=-\sum_{\mathcal{D}\in\mathcal{C}}\sum_{H_{t}\in\mathcal{D}}\log P(\mathcal{P}_{t}\mid\mathbf{E}_{t},I_{p},H_{t}),$ (7)
@@
IV-D Integration with Existing CRS Models
@@
IV-D1 Preference Representation
Traditional CRSs are not designed to directly utilize natural language preference summaries. We employ a PLM-based encoder to transform these summaries into a format suitable for existing CRS architectures. The encoder is adaptable and can be implemented as either a frozen or trainable text encoding model. We employ BERT [35] to encode the user preference summary as follows:
$\bm{s}_{c}=\text{PLM}(\mathcal{P}),$ (8)
where $\bm{s}_{c}$ is the encoded preference from text preference $\mathcal{P}$, specifically the [CLS] token embedding. The contextual understanding of language models enables $\bm{s}_{c}$ to capture comprehensive user preference summaries at both coarse-grained (i.e., overall user preferences) and fine-grained levels (i.e., current interests and specific items).
IV-D2 Enhanced Recommendation
We integrate the encoded preferences into existing CRS models to enhance their recommendation performance. We employ an adaptive gating mechanism to enhance the preference representation $\bm{s}_{b}$ captured from the base CRS model  [^1] The preference representation $\bm{s}_{b}$ is specific to each CRS model as different models employ distinct methods for building user preferences. with our COMPASS-generated representation $\bm{s}_{c}$. Note that the preference representation $\bm{s}_{b}$ is specific to each CRS model as different models employ distinct methods for building user preferences. Formally, we have:
$\gamma$ $=\sigma(\mathbf{W}[\bm{s}_{b};\bm{s}_{c}]),$ (9)
$\bm{s}_{u}$ $=\gamma\bm{s}_{b}+(1-\gamma)\bm{s}_{c},$ (10)
@@ IV-D2 Enhanced Recommendation
$P_{rec}(i)=\text{softmax}(\bm{s}_{u}\cdot\bm{I}_{i}),$ (11)
@@
$\mathcal{L}_{\text{rec}}=-\sum_{j=1}^{N}\sum_{i=1}^{M}y_{ij}\log\left(\bm{P}_{\text{rec}}^{(j)}(i)\right),$ (12)
@@
V Experiments
V-A Experimental Settings
V-A1 Datasets
TABLE I: Performance comparison on recommendation tasks. ‘COM.’ denotes models enhanced with the COMPASS approach and ‘Improv.’ indicates the relative improvement of COMPASS compared to the original base model. The best results are highlighted in bold (t-test with $p<0.05$).
@@ V-A1 Datasets
Base ReDial [4] 0.140 0.320 0.061 0.065 0.035 0.045 0.117 0.285 0.035 0.072 0.022 0.048
BERT [35] 0.143 0.319 0.073 0.108 0.052 0.059 0.179 0.328 0.095 0.125 0.072 0.079
GPT-2 [36] 0.147 0.327 0.071 0.107 0.051 0.056 0.112 0.278 0.089 0.128 0.063 0.076
Llama3.1-8B [11] 0.188 0.376 0.103 0.146 0.078 0.087 0.190 0.332 0.118 0.150 0.094 0.102
KBRD [7] 0.151 0.336 0.099 0.136 0.071 0.079 0.172 0.265 0.106 0.127 0.086 0.091
KGSF [5] 0.183 0.378 0.098 0.140 0.072 0.081 0.175 0.273 0.106 0.128 0.088 0.093
PECRS [37] 0.205 0.399 0.112 0.154 0.083 0.093 0.179 0.337 0.106 0.142 0.084 0.092
BARCOR [24] 0.169 0.374 0.088 0.133 0.063 0.073 0.185 0.339 0.104 0.137 0.080 0.087
UniCRS [8] 0.216 0.416 0.118 0.163 0.087 0.095 0.250 0.408 0.182 0.218 0.148 0.160
DCRS [38] 0.242 0.422 0.142 0.182 0.111 0.120 0.225 0.417 0.152 0.194 0.128 0.135
MSCRS [39] 0.255 0.448 0.153 0.192 0.122 0.130 0.257 0.424 0.163 0.205 0.145 0.151
@@
COM.+MSCRS 0.264 0.451 0.160 0.200 0.129 0.137 0.277 0.427 0.183 0.214 0.154 0.160
   +Improv. 3.53% 0.67% 4.58% 4.17% 5.74% 5.38% 7.78% 0.71% 12.27% 4.39% 6.21% 5.96%
We conduct our experiments on two widely used English CRS datasets: ReDial [4] and INSPIRED [40] datasets. ReDial dataset, focused on movie recommendations, contains 11,348 conversations and is constructed through crowdsourcing on Amazon Mechanical Turk (AMT). The INSPIRED dataset contains 999 conversations, also about movie recommendations, but additionally provides recommendation strategies based on social science theories. All experiments are conducted using the official dataset splits. We constructed the KG by scraping data from IMDB [^3] https://www.imdb.com/, using movie titles and release years as key search terms.
V-A2 Baselines
We consider a comprehensive range of baseline models, including traditional CRS ReDial [4], knowledge-graph based methods such as KBRD [7] and KGSF [5], language model-based methods like BERT [35], GPT-2 [36], and Llama3.1-8B [11], metadata-augmented language model PECRS [37], and hybrid approaches that combine KGs with language models: BARCOR [24] fuses BART with a KG, UniCRS [8] employs KG prompt learning, DCRS [38] leverages retrieval-augmented prompt learning, and MSCRS [39] introduces multi-modal graph prompt learning. Since no existing baselines specifically generate user preference summaries, we use Llama3.1-8B and GPT-4o as baselines for comparison, both without KG access. For clarity, we denote Llama3.1-8B as Llama-Summary when used for preference summary generation.
V-A3 Evaluation Metrics
Our evaluation framework assesses both the performance of the recommendations and the quality of generated user preference summaries. For recommendation tasks, we adopt widely used metrics from previous works [23, 5], including HR@$K$, NDCG@$K$, and MRR@$K$, with $K$ set to 10 and 50. To evaluate user preference summaries, we employ three complementary metrics: (1) Lexical Similarity: ROUGE-1, ROUGE-2, ROUGE-L, and ROUGE-Sum; (2) Semantic Understanding: Following [41], we use GPT-4o-mini to assess Reasoning Proficiency (RP) and Factual Consistency (FC). RP assesses logical coherence, inference accuracy, and recommendation relevance, while FC evaluates dialogue alignment, knowledge graph consistency, and statement accuracy. Both metrics are scored on a scale of 0 to 100. (3) User-Centric Evaluation: Inspired by [42, 43], we conduct LLM-based simulated user evaluations, which rank the system’s output based on two aspects: Explainability, which measures how clearly the system justifies its recommendations, and User Preference Alignment, which measures how accurately the system understands and responds to user preferences expressed in the conversation.
V-A4 Implementation Details
We implement COMPASS using the Llama3.1-8B model [11] as the base LLM, freeze all base LLM parameters and employ LoRA [44] for efficient fine-tuning. The graph encoder uses 1 layer with 768 hidden dimensions. For entity linking, we use provided entities for INSPIRED dataset and GPT-3.5 to extract entities from Redial dialogues, then apply text-embedding-ada-002 for semantic matching to KG entities. We use batch sizes of 256 for pre-training and 128 for fine-tuning on Inspired dataset. For ReDial dataset, we maintain a batch size of 128 for both pre-training and fine-tuning. All experiments run on Nvidia A100 GPUs. For ground truth preference summaries, we employ GPT-4 to generate summaries based on dialogue histories and associated item metadata from the KG using structured instruction templates [^4] Code, data, and instruction templates are available at: https://github.com/icedpanda/COMPASS-official.
V-B Recommendation Evaluation
V-B1 Improvement over Baseline Models
@@
COMPASS is designed to be flexible, allowing integration with different CRS models. We evaluated its effectiveness across various models, with performance results shown in Table I. We have the following observations: (1) Baseline models incorporating external KGs, such as KBRD and KGSF, consistently outperform simpler language model-based approaches like BERT and GPT-2. This highlights the importance of structured knowledge in capturing user preferences and item relationships. settings. (2) Llama3.1-8B demonstrates strong performance, surpassing BARCOR and performing competitively with PECRS on ReDial dataset, indicating that the advanced language understanding and extensive world knowledge of LLMs can effectively capture user preferences, resulting in more accurate recommendations. (3) Advanced hybrid methods that integrate language models with KGs achieve the strongest baseline performance, with MSCRS achieving the best results across all metrics on both datasets.
@@
Building upon these baselines, COMPASS consistently enhances every integrated model regardless of architecture or performance level. The framework delivers substantial gains on weaker models, achieving 44.76% HR@10 improvement with KBRD on INSPIRED dataset and 27.27% with BERT on ReDial dataset. When integrated with LLMs like Llama3.1-8B, COMPASS maintains solid enhancements of 13.76% in HR@10 on ReDial dataset. Most importantly, COMPASS enhances the strongest baseline MSCRS, achieving 7.78% HR@10 improvement and reaching the best overall performance across all metrics. These consistent improvements across different architectures validate that COMPASS’s knowledge-enriched preference representations provide universal value. These representations effectively capture user intent while providing structured insights that enhance both recommendation performance and explainability.
V-B2 Comparison of Enhancement Methods